Continue the two sequences of numbers below and find an equation to each of the sequences:
n 1 2 3 4 5 6 7 Equation
an 2 5 9 14 20 27
bn 1 3 12 60 360 2520
1
Expert's answer
2021-10-19T15:59:02-0400
"a_n=a_{n-1}+n+1, n\geq2, a_1=1"

"a_2=a_1+2+1=2+3=5"

"a_3=a_2+3+1=5+4=9"

"a_4=a_3+4+1=9+5=14"

"a_5=a_4+5+1=14+6=20"

"a_6=a_5+6+1=20+7=27"

"a_7=a_6+7+1=5+4=27+8=35"




"b_n=b_{n-1}(n+1), n\geq2, b_1=1"

"b_2=b_1(2+1)=3"

"b_3=b_2(3+1)=12"

"b_4=b_3(4+1)=60"

"b_5=b_4(5+1)=360"

"b_6=b_5(6+1)=2520"

"b_7=b_6(7+1)=20160"